Zn + HCI > ZnCI2 + H2

A. What is my theoretical yield of H2 if I start with 74 g of HCI?

B. If my percent yield was 93%, what was my actual yield?
1
Expert's answer
2020-02-28T12:14:15-0500

Solution.

1)

"Zn + 2HCl = ZnCl2 + H2"

"n(HCl) = \frac{m(HCl)}{M(HCl)}"

n(HCl) = 2.03 mol

"n(H2) = \frac{2.03}{2} = 1.015 \ mol"

"V(H2) = n(H2) \times Vm = 22.74 \ L"

2)

If we consider the percentage yield of zinc chloride, we find the theoretical mass of zinc chloride.

"n(ZnCl2) = \frac{2.03}{2} = 1.015 \ mol"

m(ZnCl2) = 138.04 g

"m(pract.) = y \times m(theory)"

m(pract.) = 128.38 g

Answer:

V(H2) = 22.74 L

m(pract.) = 128.38 g